OpenStudy (anonymous):

function f(x) is reflected in x axis then shifted up 4 units, and the graph of g(x) = 4-x^2 is obtained. what is f(x)?

OpenStudy (amistre64):

in order to shift a graph "up"; we have to add a number to it; to determine the original place for it; subtract that number back out.

OpenStudy (amistre64):

4-x^2 -4 ------- -x^2 = f(x)

OpenStudy (amistre64):

ack .. for got to reflect it :)

OpenStudy (amistre64):

change the sign and you got it

OpenStudy (anonymous):

look 1 up above plz

OpenStudy (anonymous):

so wait what do i get for answer?

OpenStudy (anonymous):

what is ack?

OpenStudy (amistre64):

"ack" just means that I left something out that I forgot to include

OpenStudy (anonymous):

amistre64 look 1 up above plz

OpenStudy (amistre64):

f(x) reflects over the x axis means it turns to -f(x) add 4 means we get: -f(x) + 4 -f(x) + 4 = 4-x^2 ; we can solve for f(x) now, subtract 4 -f(x) = -x^2 ; multiply by -1 f(x) = x^2